the frequency distribution below summarizes employee years of service for a regional hospital. determine the width of each class.
years of service   / frequency
1-5                          /   5
6-10                        / 20
11-15                        /25
16-20                         / 10
21-25                           /5
26-30                             / 3
(a) 10 years, ( b) 5 years, (c) 4 years, (d) 6 years
Answer #1

Answer)

Width is given by

Upper class limit - lower class limit

= 5-1 = 10-6 = 15-11 = 20-16 = 25-21 = 30-26 = 4 years

Answer is option C